OpenStudy (anonymous):

Help i got to help my little sister with her math and i have no idea how to do this!! Solve the quadratic equation (2x-1)(x+7)=0

OpenStudy (anonymous):

it would be easier to help your sister she probably just needs a little more than she knows but your question just wants you to find a value for x 2x-1=0 is one part so x=1/2 would work and second part x+7 =0 x=-7 would work best of luck

OpenStudy (anonymous):

do distribute property